OpenStudy (anonymous):

is there a simpler form of 5/(tan(x)+sec(x)) ?

OpenStudy (jdoe0001):

\(\bf \cfrac{5}{tan(x)}+sec(x)?\)

OpenStudy (anonymous):

\[5/(tanx+secx)\]

OpenStudy (jdoe0001):

hmm \(\bf \cfrac{5}{tan(x)+sec(x)}\implies \cfrac{5}{\frac{sin(x)}{cos(x)}+\frac{1}{cos(x)}}\implies \cfrac{5}{\frac{sin(x)+1}{cos(x)}} \\ \quad \\ \cfrac{\frac{5}{1}}{\frac{sin(x)+1}{cos(x)}}\implies \cfrac{5}{1}\cdot \cfrac{cos(x)}{sin(x)+1}\implies\cfrac{5cos(x)}{sin(x)+1}\) not sure if I'd caller a simplified version of it though

OpenStudy (anonymous):

haha- yeah thats where i called it quits and came here